Apex Fitness Club uses straight-line depreciation for a machine costing $26,850, with an estimated four-year life and a $2,900 salvage value. At the beginning of the third year, Apex determines that the machine has three more years of remaining useful life, after which it will have an estimated $2,450 salvage value.

Required:
1. Compute the machine’s book value at the end of its second year.
2. Compute the amount of depreciation for each of the final three years given the revised estimates.
Answer #1

Depreciation = (Cost - Salvage value) /Useful life

= (26,850-2900)/4 = 8987.50 per year

1. Book value = 26,850 - (5987.50*2)

= 14,875

2. Depreciation = (14,875 - 2,450)/3

= 4142

I have rounded some numbers. Comment if you face any issues. I will edit answer immediately